# Magnetoconvection in a spherical shell: Equatorial symmetry during the transition from the weak- to the strong-field regime

### 09 Mar 2026 v1

@@ Appendix B Equatorial symmetry
Roberts (1968) derived the equations governing the onset of hydrodynamic convection in a sphere via internal heat sources; he noted that, after considering solutions in the form of spherical harmonics, the structure of the resultant equations leads to two distinct families of solutions emerging, which are characterised by the parity of $l+m$, where $l,m$ are the spherical harmonic degree and order, respectively, of $U$ and $T$, where
@@
In the axisymmetric, $m=0$, case, Roberts (1968) shows that the odd modes are the most unstable - this solution corresponds to a global meridional circulation, with streamlines corresponding to hemispherical rings. Roberts similarly calculates the critical Rayleigh numbers for odd modes with $m>0$. For this reason, odd modes can be referred to as Roberts symmetry.
Busse (1970) then tackles a perturbation analysis for the spherical convection problem and calculates the critical Rayleigh numbers of the even modes, showing that these are preferred at onset for $m>0$ and generally preferred when considering all $m$, for small Ekman numbers. These modes include tall thin columnar structures appearing at onset as thermal Rossby waves. This symmetry is sometimes called Busse symmetry and is frequently taken as an assumption in linear onset convection studies to simplify the analysis.
@@
If an inner core is present, global meridional convection becomes impossible, and large scale meridional convection must be restricted to the outside of the tangent cylinder (Figure 20a). Smaller convection cells could also develop inside the tangent cylinder (Figure 20b); these latter modes can be efficient in transporting heat, since streamlines abut both the inner and outer boundaries, however, they are impeded by viscosity due to their small scale. Gilman (1975) refers to these modes as polar convective modes. In the absence of the region outside the tangent cylinder the two polar regions would be disconnected and there would be no preference for either equatorially symmetric or antisymmetric modes; if we consider the outer continuation of this mode, then the Roberts mode, which aligns with the meridional circulation in Figure 20a, will be slightly preferable at onset (Sakuraba, 2002).
@@
Fearn (1979) shows that a toroidal field can destabilise the system so that convection occurs at a lower critical Rayleigh number, provided the magnetic field strength is large enough. The most unstable modes in this case are modes with Busse symmetry which are either modified Rossby modes or magnetostrophic modes when $\Lambda$ is sufficiently large. Sakuraba (2002) investigates the effect of an axial magnetic field and shows that the polar convective mode can be destabilised at much smaller $\Lambda$ (in fact, Sakuraba finds that an antisymmetric mode can sometimes be the preferred mode of convection).

### 17 Mar 2026 v2

@@ Appendix B Equatorial symmetry
Roberts Roberts (1968) derived the equations governing the onset of hydrodynamic convection in a sphere via internal heat sources; he noted that, after considering solutions in the form of spherical harmonics, the structure of the resultant equations leads to two distinct families of solutions emerging, which are characterised by the parity of $l+m$, where $l,m$ are the spherical harmonic degree and order, respectively, of $U$ and $T$, where
@@
In the axisymmetric, $m=0$, case, Roberts (1968) show that the odd modes are the most unstable - this solution corresponds to a global meridional circulation, with streamlines corresponding to hemispherical rings. Roberts similarly calculates the critical Rayleigh numbers for odd modes with $m>0$. For this reason, odd modes can be referred to as Roberts symmetry.
Busse Busse (1970) then tackles a perturbation analysis for the spherical convection problem and calculates the critical Rayleigh numbers of the even modes, showing that these are preferred at onset for $m>0$ and generally preferred when considering all $m$, for small Ekman numbers. These modes include tall thin columnar structures appearing at onset as thermal Rossby waves. This symmetry is sometimes called Busse symmetry and is frequently taken as an assumption in linear onset convection studies to simplify the analysis.
@@
If an inner core is present, global meridional convection becomes impossible, and large scale meridional convection must be restricted to the outside of the tangent cylinder (Figure 20a). Smaller convection cells could also develop inside the tangent cylinder (Figure 20b); these latter modes can be efficient in transporting heat, since streamlines abut both the inner and outer boundaries, however, they are impeded by viscosity due to their small scale. Gilman Gilman (1975) refers to these modes as polar convective modes. In the absence of the region outside the tangent cylinder the two polar regions would be disconnected and there would be no preference for either equatorially symmetric or antisymmetric modes; if we consider the outer continuation of this mode, then the Roberts mode, which aligns with the meridional circulation in Figure 20a, will be slightly preferable at onset (Sakuraba, 2002).
@@
Fearn Fearn (1979) shows that a toroidal field can destabilise the system so that convection occurs at a lower critical Rayleigh number, provided the magnetic field strength is large enough. The most unstable modes in this case are modes with Busse symmetry which are either modified Rossby modes or magnetostrophic modes when $\Lambda$ is sufficiently large. Sakuraba Sakuraba (2002) investigates the effect of an axial magnetic field and shows that the polar convective mode can be destabilised at much smaller $\Lambda$ (in fact, Sakuraba finds that an antisymmetric mode can sometimes be the preferred mode of convection).